# Piezoelectric actuation for integrated photonics

### v1

@@ 8 Conclusion
Equipped with DC tuning and AOM, the concept of a multifunctional photonic processor is conceivable, as depicted in Fig. 1 at the outset of this review. Ranging from tunable laser sources and programmable beam splitters to optical modulators, the Piezo-on-Photonic platform facilitates the integration of these elements on a single chip, albeit requiring moderate adjustments in device design and fabrication processes to suit various materials, etching selectivity, and release compatibility. Looking ahead, piezoelectric actuators are poised to significantly contribute to the development of programmable photonic circuits across spatial and frequency domains, enabling sophisticated optical simulation and computation. Furthermore, investigating new piezoelectric materials like AlScN [453, 454], BaTiO3 [121], and HZO [455] could enhance tuning and modulation efficiency, thus reducing the required drive voltage and shrinking the device size. Additionally, exploring innovative fabrication techniques for the monolithic integration of III-V gain components on the same piezoelectric-actuated photonic chip could provide tunable lasers with reduced noise and enable large-scale production [246, 456]. Consequently, we anticipate that piezoelectric actuation will evolve into an essential component in future photonic circuit designs, offering distinctive tuning and modulation capabilities with high efficiency, minimal power requirements, and cost-effectiveness.
9 Acknowledgement
@@
We thank Scott E. Kenning and Jiahao Sun for valuable discussion and critical reading of this review. We thank Bin Dong, Michael Zervas, Anat Siddharth, Rui Ning Wang, Terence Blésin, Jijun He, Rui Ning Wang, Erwan Lucas, Arslan S. Raja, Grigory Lihachev, Tianyi Liu, Miles H. Anderson, Wenle Weng, Ben Yu, Mert Torunbalci, Alaina Attanasio, Wil Kao, Andrea Bancora, Viacheslav Snigirev, Johann Riemensberger, Vladimir Shadymov, Diego Visani, Andrey Voloshin for contributions to the fabrication and experimental realization of piezoelectric actuated photonic devices. We also thank the collaborations from Christian Koos, Yung Chen, and Huanfa Peng. We acknowledge the support from the Defense Advanced Research Projects Agency (DARPA) Microsystems Technology Office (MTO) under contract HR0011-15-C-055 (DODOS), the Air Force Office of Scientific Research under award no. FA9550-19-1-0250 and no. FA8655-20-1-7009, US National Science Foundation’s RAISE TAQS program under grant no. PHY 18-39164, the EU H2020 research and innovation program under grant agreement no. 732894 (HOT), the Swiss National Science Foundation under grant agreement no. 176563 (BRIDGE), and the Marie Sklodowska-Curie IF grant agreement No. 101033663 (RaMSoM). H.T. acknowledges the support from NSF QISE-Net under grant no. DMR 17-47426. We thank the support from the EPFL centre of MicroNanoTechnology (CMi) and the Birck Nanotechnology Center at Purdue University for the fabrication of integrated Si3N4 photonic chips and AlN actuators. We also thank Plasma-Therm LLC. and Advanced Modular Systems Inc. for AlN deposition, and Radiant Technologies Inc. for fabricating PZT actuators.

### v2

@@ 8 Conclusion
Equipped with DC tuning and AOM, the concept of a multifunctional photonic processor is conceivable, as depicted in Fig. 1 at the outset of this review. Ranging from tunable laser sources and programmable beam splitters to optical modulators, the Piezo-on-Photonic platform facilitates the integration of these elements on a single chip, albeit requiring moderate adjustments in device design and fabrication processes to suit various materials, etching selectivity, and release compatibility. Looking ahead, piezoelectric actuators are poised to significantly contribute to the development of programmable photonic circuits across spatial and frequency domains, enabling sophisticated optical simulation and computation.
In the future, advances in nanofabrication technologies, driven by breakthroughs in chemical and material science, could enhance the performance of piezoelectric actuators and the creation of cutting-edge functional devices. For instance, the epitaxial growth of single crystalline piezoelectric thin films at the wafer scale that is CMOS compatible, such as GaN [509] and AlN [510], will attract more research interest due to their lower acoustic and optical losses, negligible leakage current, and better piezoelectric response. Furthermore, investigating novel piezoelectric materials with higher piezoelectric coefficient, such as AlScN [128], BTO [110], HZO [246, 477], and KNN [478], could enhance the tuning and modulation efficiency, and thus reduce the required drive voltage and shrink the device size.
Finally, exploring innovative fabrication techniques for the monolithic integration of III-V gain components on the same piezoelectric-actuated photonic chip could provide tunable lasers with reduced noise and enable large-scale production [259, 511]. Consequently, we anticipate that piezoelectric actuation will evolve into an essential component in future photonic circuit designs, offering distinctive tuning and modulation capabilities with high efficiency, minimal power requirements, and cost-effectiveness.
Funding
U.S. National Science Foundation’s RAISE TAQS program (PHY 18-39164); Air Force Office of Scientific Research (FA8655-20-1-7009); EU H2020 research and innovation programme (732894 HOT); Swiss National Science Foundation (SNSF) (176563 and 211728 BRIDGE); Defense Advanced Research Projects Agency (DARPA) (W911NF212024 NINJA LASER); Microsystems Technology Office (MTO) (HR0011-15-C-055 DODOS).
Acknowledgement
@@
We thank Jiahao Sun for valuable discussion and critical reading of this review. We thank Bin Dong, Michael Zervas, Erwan Lucas, Arslan S. Raja, Tianyi Liu, Miles H. Anderson, Wenle Weng, Ben Yu, Mert Torunbalci, Wil Kao, Diego Visani, Jijun He for valuable discussion during the fabrication and experimental realization of piezoelectric actuated photonic devices. We also thank the collaborations from Christian Koos, Yung Chen, and Huanfa Peng. We acknowledge the support from the Defense Advanced Research Projects Agency (DARPA), Microsystems Technology Office (MTO), Air Force Office of Scientific Research, U.S. National Science Foundation, the EU H2020 research and innovation program, and the Swiss National Science Foundation (SNSF). H.T. acknowledges the support from NSF QISE-Net under grant no. DMR 17-47426. We thank the support from the EPFL centre of MicroNanoTechnology (CMi) and the Birck Nanotechnology Center at Purdue University for the fabrication of integrated Si3N4 photonic chips and AlN actuators. We also thank Plasma-Therm LLC. and Advanced Modular Systems Inc. for AlN deposition, and Radiant Technologies Inc. for fabricating PZT actuators.
Disclosures
TJK: LiGenTec SA (I, P), DEEPLIGHT SA (I, P); AB: DEEPLIGHT SA (I, E, P); AV: DEEPLIGHT SA (I, P)
Data Availability
No data were generated or analyzed in the presented research.